OpenStudy (secret-ninja):

Am I right? Math help. Medals. What is the solution to the equation 2(3)^x = 3^x + 1?

OpenStudy (secret-ninja):

@sidsiddhartha

OpenStudy (secret-ninja):

Choices: x = -1 x = 0 x = 1 x = 2

OpenStudy (secret-ninja):

I think its 0.

OpenStudy (anonymous):

Yups. 2(3)^0 = 2(1) 3^0 + 1 = 2